Define the normalized frequency for an optical fiber and explain its use in the determination of the number of guided modes propagating within a step index fiber. A step index fiber in air has a numerical aperture of 0.16, a core refractive index of 1.45 and a core diameter of 60 μm. Determine the normalized frequency for the fiber when light at a wavelength of 0.9 μm is transmitted. Further, estimate the number of guided modes propagating in the fiber.


1
Expert's answer
2021-04-13T06:32:11-0400

The normalized frequency for an optical fiber is defined as

"V = \frac{2πa}{λ}NA"

NA – numerical aperture

a – radius of the core

λ – wavelength

Normalized frequency is used to calculate the number of modes M:

"M = \frac{V^2}{2}"

Given:

NA = 0.16

Core diameter = 60 μm

Core radius a = 30 μm

λ = 0.9 μm

"V = \frac{2π \times 30 \times 10^{-6} \times 0.16}{0.9 \times 10^{-6}} = 33.51 \\

M = \frac{33.51^2}{2}=561.47"